a fan rotating 100 rad/sec is switched off after 2n rotation its angular velocity becomes 50 rad/sec find its angular velocity after n revolution

a fan rotating  100 rad/sec is switched off after 2n rotation its angular velocity becomes 50 rad/sec find its angular velocity after n revolution

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Angular displacement in 2n rotations is4n\pi.
Hence from equation of motion,

\omega^2-\omega_0^2=2\theta \alpha
\implies \alpha=\dfrac{50^2-100^2}{2(2n\pi)}=-\dfrac{7500}{8n\pi}
Hence angular velocity after n rotations=\sqrt{\omega_0^2+2\alpha\theta_2}=\sqrt{100^2-2\dfrac{7500}{8n\pi}(2n\pi)}=79rad/sec